8. Calculate the pH of a 0.00175 M solution of KOH. 9. Calculate the (H+] in a solution that has a pH of 8.38.
Answer #1

Col 0.00175 M pot z log [on] = -log(0-00175) pon = 2.757 pl a plew-bot = 14-2.757 = 11,243 9 pl 28.38 [at] z 10-pH = 10-8.38